OpenStudy (dtan5457):

How do you find asymptotes on a tangent graph? Like 2tan3x. I have the graph labeled, I have the period, amplitude, etc, I just don't know how to get the asymptotes.

OpenStudy (dtan5457):

@Nnesha @freckles @jim_thompson5910

OpenStudy (dtan5457):

@SolomonZelman

OpenStudy (loser66):

ok, look at tan, tan= sin/cos, right? hence, it is undefined when cos =0

OpenStudy (loser66):

when cos =0? it happens if x = pi/2, 3pi/2, right?

OpenStudy (loser66):

dat sit. or if you remember, tan function is undefined if x = those guys.

OpenStudy (dtan5457):

Why does tan= sin/cos? Is there a formula to remember here?

OpenStudy (loser66):

I think so. my prof told me that tan = sin/cos

OpenStudy (freckles):

 \[\sin(\theta)=\frac{y}{r} \\ \cos(\theta)=\frac{x}{r} \\ \tan(\theta)=\frac{y}{x} \text{ which is another way to express } \frac{\frac{y}{r}}{\frac{x}{r}} \\ \text{ which is another way to express } \frac{\sin(\theta)}{\cos(\theta)}\]